[Hallc_running] [New Logentry] RC Daily Update

pooser at jlab.org pooser at jlab.org
Sat Nov 17 10:05:01 EST 2018

Logentry Text:
RC report:
We began yesterday day shift with collecting data for kinematic group 9-9.  Beam was stable throughout day shift (https://logbooks.jlab.org/entry/3627651) and we were able to complete 4 kinematic subgroups and the transition into swing shift was smooth.  

At 18:01 Dave G. Noticed that the xscaler GUI was not updating in the usual way (https://logbooks.jlab.org/entry/3627385).  In order to recover the scaler readout, all ROCs were rebooted (https://logbooks.jlab.org/entry/3627394) and CODA was rebooted (https://logbooks.jlab.org/entry/3627392) however, CODA was not able to connect to VX works ROC’s.  Multiple reboots were attempted to no avail and Brad S. was contacted to debug the issue. He discovered that the issue was related to the OOM killer on cdaql3 clobbering NFS (hence CODA) related processes (https://logbooks.jlab.org/entry/3627420).  Data collection resumed promptly after the Brad’s fix was implemented and group 9-10 was completed and 9-11 commenced during the remainder of swing shift.

During owl run 6277 CODA crashed and was rebooted, then during run 6278 it was noticed that 4 sets (discriminator cards) of 16 HMS DC wires were missing (https://logbooks.jlab.org/entry/3627562, https://logbooks.jlab.org/entry/3627583).  ROC3 was then rebooted and the missing channels were recovered with run 6281 (https://logbooks.jlab.org/entry/3627588).  Data collection was smooth throughout the remainder of the night and kinematic group 9 was completed at 07:40 at which point the spectrometers were configured for kinematic setting 13.

At the beginning of day shift, while attempting to rotate the spectrometers according to kinematic group 13, an emergency stop was initiated via the HMS encoder (https://logbooks.jlab.org/entry/3627674) which in turn tripped the shunt breakers.  An access was made to reset the breakers (https://logbooks.jlab.org/entry/3627693) and the spectrometers were successfully rotated to the desired angles (https://logbooks.jlab.org/entry/3627695).  Data collection for kinematic group 13 commenced at 09:30 which is where we will likely remain for the next day or so if beam remains reliable.

Notices in effect for shift crews:
Shift leaders, please fill out the check list once per shift!
It is no longer required to fill out the standard.kinematics file online, it is now done automagically
Split the large desired charge between multiple runs (~1 hour per run)
Listed currents are estimates. For runs at less than 60uA, increase current until SHMS singles rates are between 400k-500k.
How-to power cycle PLC I/O chassis: https://logbooks.jlab.org/entry/3625862, https://logbooks.jlab.org/entry/3625900

Planned accelerator configuration changes:
Tuesday, Nov 20: CHL UPS battery change, should be noninvasive however, induce CHL anomaly is possible
Wednesday, Nov 21: Hall A goes offline and MCC wants to conduct high current beam study of energy spread at 40 uA.  It is unclear that we will be effected
Monday, Nov 26: Energy change to 1.86 GeV/pass, we will be moved over to 500 MHz laser.

Planned accesses to the hall:
Monday, Nov 26: aerogel change and HGC work needs to be coordinated with Walter K.

Special requests:
Vijay Kumar (Regina group) has requested a walkthrough during the next access
Jay B. would like to conduct another beam steering test (5 minute procedure) during the next polarity change.  He requests to only be contacted if the polarity change occurs during the hours of 0630 and 1930.
Carlos Munoz Camacho would like access to the hall to show Hall-C to two French engineers  working on the NPS project.  It should be a ~30 minute access, and is a low priority.  He can be contacted at 357-1065

There will be no RC meetings this weekend, the next meeting will take place at 16:30 on Monday 11/19.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3627722
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181117/64ebb14b/attachment.html>

More information about the Hallc_running mailing list